OpenStudy (anonymous):

hi, I have to find the short leg, long leg, and hypotenuse for a right triangle. The short leg = x, the long leg = 1/2x+11 and the hypotenuse is 2x+1. Using a^2+b^2=c^2, I got x^2+(1/2x+11)^2=2x+1^2 but now I'm stumped and I'm not sure what to do next... please help

jimthompson5910 (jim_thompson5910):

you should have this x^2+(1/2x+11)^2=(2x+1)^2 notice the (2x+1)^2

jimthompson5910 (jim_thompson5910):

since you're squaring all of 2x+1

OpenStudy (anonymous):

ah, sorry, thank you for fixing that! so, x^2+(1/2x+11)^2=(2x+1)^2. my dad and I are trying to solve this but we're not sure how to multiply now

jimthompson5910 (jim_thompson5910):

you now need to FOIL out (1/2x+11)^2 and (2x+1)^2

OpenStudy (anonymous):

Okay we got 1/4x^2+11x+121 and 4x^2+2x+1

jimthompson5910 (jim_thompson5910):

(2x+1)^2 foils out to 4x^2 + 4x + 1 the other one looks good

OpenStudy (anonymous):

okay you're right thank you for fixing that

jimthompson5910 (jim_thompson5910):

now get everything to one side and simplify/combine like terms after that, you use the quadratic formula

OpenStudy (anonymous):

We're not really sure if we combined it right we got 11/4x^2-7x-120=0

jimthompson5910 (jim_thompson5910):

you are correct, now solve that using the quadratic formula

jimthompson5910 (jim_thompson5910):

Optional: Multiply both sides by 4 to clear out the fraction 11/4x^2-7x-120=0 4*(11/4x^2-7x-120)=4*0 11x^2 - 28x - 480 = 0 then use the quadratic formula

OpenStudy (anonymous):

okay so plugging that into the quadratic formula we got x=176 and x=-120 and it can't be negative, so the answer for x would be 176?

jimthompson5910 (jim_thompson5910):

I'm not getting either of those possible solutions. I'm getting two completely different numbers. So try again

OpenStudy (anonymous):

ahh! i forgot to divide it T.T okay, i divided it this time and i got x=8

jimthompson5910 (jim_thompson5910):

yep, the final answer is x = 8

OpenStudy (anonymous):

oh my gosh ;-; thank you so so so so much!!!

jimthompson5910 (jim_thompson5910):

you're welcome